OpenStudy (anonymous):

find the integral : {cos^3x+5/cos^2x dx

OpenStudy (zarkon):

this? \[\int\frac{\cos^3(x)+5}{\cos^2(x)}dx\]

OpenStudy (anonymous):

؟

OpenStudy (anonymous):